In one camera, the resistance of the bulb is 0.1 Ω, and the capacitor has a value of 4700 μF. How long will it take to discharge the capacitor to just 5% of its original voltage?


1
Expert's answer
2021-01-28T18:12:05-0500

Let's first find the time constant:


"\tau=RC=0.1\ \Omega\cdot4700\cdot10^{-6}\ F=4.7\cdot10^{-4}\ s."

Let's write the equation for voltage across the capacitor:


"V=V_0e^{-\dfrac{t}{\tau}},""0.05V_0=V_0e^{-\dfrac{t}{\tau}},""ln(0.05)=ln(e^{-\dfrac{t}{\tau}}),""-\dfrac{t}{\tau}=ln(0.05),""t=-\tau ln(0.05)=-4.7\cdot10^{-4}\cdot ln(0.05)=1.4\cdot10^{-3}\ s."